OpenStudy (anonymous):

which of the following points are solutions to the inequality below. y >= -3x+2 a (2,1) b 2,7) c(-1 ,10) d(0,2) e (1,-2) f(0,-1)

OpenStudy (anonymous):

pretend the > is a = sign...

OpenStudy (anonymous):

blahhhh >,<

OpenStudy (anonymous):

\[\ge \]

OpenStudy (anonymous):

@phi might can help you... lol

OpenStudy (anonymous):

c: id love them forevers

OpenStudy (anonymous):

@hartnn could you help

hartnn (hartnn):

plug in points one by one to see whether left side = right side or not, i'll do the 1st one for you,

hartnn (hartnn):

y >= -3x+2 2,1 so, x= 2, y=1 1 >= -3(2)+2 1> = -6+2 1>= -4 this is TRUE. so, 2,1 is a solution!

hartnn (hartnn):

do same thing for all other points

OpenStudy (anonymous):

okie dokie & thanks c: